OpenStudy (anonymous):

Find the most general antiderivative of the function. Use C for the constant of the antiderivative.

OpenStudy (anonymous):

\[f(x)=9\sqrt{x}+4\cos x\]

OpenStudy (dumbcow):

sqrt x = x^1/2 use power rule to find anti derivative \[\int\limits x^{n} = \frac{x^{n+1}}{n+1}\] antiderivative of cos x is sin x

OpenStudy (anonymous):

im getting 6^3/2 -sinx

OpenStudy (anonymous):

does this sound correct?

OpenStudy (dumbcow):

first part is...but you dropped the 4 with the sin x \[\int\limits 4\cos x = 4\sin x\]

OpenStudy (anonymous):

got it so it would be 6^3/2-4sinx that makes sense

OpenStudy (dumbcow):

why is it minus 4sin x ?

OpenStudy (anonymous):

for some reason i thought the antideriv of cos was (-sin) is is positive sin right?

OpenStudy (dumbcow):

f(x) ---> f'(x) sin cos cos -sin so antiderivative of cos is positive sin

OpenStudy (anonymous):

thank you very much for your time and effort

OpenStudy (dumbcow):

yw